# Orbifold modifications of complex analytic varieties

### v4

@@ 3.A. Functorial resolution of singularities
Let $X$ be a complex analytic variety. Then $X$ admits a functorial (or canonical) desingularization $r\colon\widetilde{X}\to X$. More precisely, $\widetilde{X}$ is a smooth complex analytic variety, $r$ is a projective bimeromorphic morphism, which is a composition of blowups at smooth centers contained in $X_{\operatorname{\textsubscript{\rm sing}}}$, the $r$-exceptional locus is a snc divisor, and the construction of $r$ commutes with local analytic isomorphisms.
@@
Let $X$ be a complex analytic variety, and let $\mathcal{I}$ be a coherent ideal sheaf on $X$. Then $\mathcal{I}$ admits a functorial (or canonical) principalization $r\colon\widetilde{X}\to X$. More precisely, $\widetilde{X}$ is a smooth complex analytic variety, $r$ is a projective bimeromorphic morphism, which is a composition of blowups at smooth centers contained in the cosupport of $\mathcal{I}$, $r^{-1}\mathcal{I}\cdot\mathcal{O}_{\widetilde{X}}$ is an invertible sheaf which defines a divisor of snc support, and the construction of $r$ commutes with local analytic isomorphisms.
We will need the following two results on partial desingularizations, which allow us to work locally around components of the singular locus. For example, we assume that $X$ is a complex analytic variety, and that $S_{1}$ and $S_{2}$ are disjoint compact irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$. Let $X_{1}$ and $X_{2}$ be disjoint open subsets of $X$, which contain $S_{1}$ and $S_{2}$ respectively. Suppose that, for $i=1,2$, we have a proper bimeromorphic modification $f_{i}\colon Y_{i}\to X_{i}$, which is an isomorphism over the smooth locus of $X_{i}$. We remark that there can be an obstruction to glue the modifications $f_{i}$ into a modification on $X$. Indeed, there may be a component $S_{3}$ of $X_{\operatorname{\textsubscript{\rm sing}}}$ which intersects both $S_{1}$ and $S_{2}$. If the modification of $f_{1}$ at general points of $S_{3}\cap X_{1}$ is different from the one of $f_{2}$ at general points of $S_{3}\cap X_{2}$, then we cannot glue $f_{1}$ and $f_{2}$. With the following two results, we can first obtain a modification $\rho\colon X^{\prime}\to X$ so that the singular loci of $\rho^{-1}(X_{1})$ and $\rho^{-1}(X_{2})$ are compact subsets. Then the modifications on $\rho^{-1}(X_{1})$ and $\rho^{-1}(X_{2})$, which are isomorphic over the smooth loci, are independent and we can always glue them by using Lemma 2.3.
@@
Let $X$ be a normal complex analytic variety and let $S$ be the union of some irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$. Let $\Delta$ be a reduced divisor in $X$, which does not contain any component of $S$. Then there is a projective bimeromorphic morphism $f\colon Y\to X$ satisfying the following properties. Let $U=X\setminus S$ and $V=f^{-1}(U)$.
@@
Let $r\colon Z\to X$ be a log resolution of $(X,\Delta)$ such that it can be decomposed into a sequence
$Z=Z_{m}\to\cdots\to Z_{0}=X,$
where each $Z_{j+1}\to Z_{j}$ is the blowup at a smooth irreducible center $C_{j}$ contained in the singular locus (i.e. the non-snc locus) of $(X_{j},\Delta_{j})$, and $\Delta_{j}$ is the sum of the strict transform of $\Delta$ in $X_{j}$ and the whole reduced exceptional divisor of $X_{j}\to X$. We will construct by induction a sequence of projective bimeromorphic morphisms
$Y=Y_{m}\to\cdots\to Y_{0}=X$
such that the following properties hold. The composition $f\colon Y\to X$ satisfies the properties of the proposition. In particular, we can define $S_{j}$ the strict transform of $S$ in $Y_{j}$, and $Y_{j+1}\to Y_{j}$ is an isomorphism over general points of every irreducible component of $S_{j}$. Let $V_{j}$ be the preimage of $U$ in $Y_{j}$ and let $W_{j}$ be the preimage of $U$ in $Z_{j}$. Then $V_{j}\cong W_{j}$ under the natural bimeromorphic map $\psi_{j}\colon Z_{j}\dashrightarrow Y_{j}$ over $X$.
Assume that we have constructed $Y_{j}$. If $C_{j}\cap W_{j}=\emptyset$, then we let $Y_{j+1}\to Y_{j}$ be the identity map. We note that $W_{j+1}\cong W_{j}$. Hence $V_{j+1}\cong W_{j+1}$ under $\psi_{j+1}$.
If $C_{j}\cap W_{j}\neq\emptyset$, then we can define the strict transform $D_{j}$ of $C_{j}$ via $\psi_{j}$. Let $Y_{j+1}\to Y_{j}$ the blowup of $D_{j}$. It follows that $V_{j+1}\cong W_{j+1}$ under $\psi_{j+1}$. It remains to prove that $Y_{j+1}\to Y_{j}$ is an isomorphism over general points of every irreducible component of $S_{j}$. It is equivalent to show that $D_{j}$ does not contain any of these irreducible components. Assume by contradiction that $D_{j}$ contains a component $F_{j}$ of $S_{j}$. Then the image $D$ of $D_{j}$ in $X$ contains the one $F$ of $F_{j}$. We note that $D\cap U\neq\emptyset$. On the one hand, by assumption, $F$ is an irreducible component of $S$. On the other hand, $D$ is also the image of $C_{j}$ in $X$. By assumption on $Z$, $D$ is contained in singular locus of $(X,\Delta)$. It follows that there is an irreducible component $P$ of the singular locus of $(X,\Delta)$ such that
@@
In conclusion, $Y_{j+1}\to Y_{j}$ is an isomorphism over general points of every irreducible component of $S_{j}$. This completes the proof of the proposition. ∎
@@
With the notation in Proposition 3.3, we assume that $X$ has quotient singularities around $S\setminus\Delta$. Let $g\colon X^{\prime}\to Y$ be the normalization of the blowup of $Y$ at the ideal sheaf $\mathcal{I}$ of $E+f^{-1}_{*}\Delta$, and let $\rho\colon X^{\prime}\to X$ be the natural morphism. Then the following properties hold.
@@ 3.B. Blowups at an ideal sheaf
Let $X$ be a complex analytic space and let $S\subseteq X$ be a closed subspace. If $g\colon S^{\prime}\to S$ is a blowup at a coherent ideal sheaf $\mathcal{I}$, then there is a natural blowup $f\colon X^{\prime}\to X$, such that the restriction of $f$ on the strict transform of $S$ in $X^{\prime}$ is isomorphic to $g$. Indeed, the closed analytic subspace $Z\subseteq S$ defined by $\mathcal{I}$ is naturally a closed analytic subspace of $X$. If we denote the ideal sheaf of $Z$ in $X$ by $\mathcal{I}_{1}$, then we set $f\colon X^{\prime}\to X$ as the blowup at the ideal sheaf $\mathcal{I}_{1}$. As a consequence, if $g\colon S^{\prime}\to S$ is now a composition of a sequence of functorial blowups, then there is a natural functorial morphism $f\colon X^{\prime}\to X$ such that the restriction of $f$ on the strict transform of $S$ in $X^{\prime}$ is isomorphic to $g$.
For an effective Cartier divisor $H\subseteq S$, the blowup of $S$ at (the ideal sheaf of) $H$ is just an isomorphism. However, we can blowup $X$ at $H$ and get a non trivial morphism $f\colon X^{\prime}\to X$. In this case, $f$ induces an isomorphism from the strict transform $S^{\prime}$ of $S$ to $S$. We will focus on this operation in the next subsection.
In the remainder of this subsection, we let $o\in S\subseteq X$ be a point. We will shrink $S$ and $X$ around $o$ freely. Let $W$ be an open neighborhood of $\{\mathbf{0}_{N}\}\times S$ in $\mathbb{C}^{N}\times S$. We assume that there is a closed immersion from $X$ to $W$ such that the composite inclusion $S\subseteq X\subseteq W$ coincides with $\{\mathbf{0}_{N}\}\times S\subseteq W$. We denote by $\mathcal{J}_{X}$ the ideal sheaf of $X$ inside $W$. Let $F(x_{1},...,x_{N},s)$ be a holomorphic function on $W$, where $(x_{1},...,x_{N})$ are the coordinates of $\mathbb{C}^{N}$ and $s\in S$ is a point. We assume that $F\in\mathcal{J}_{X}$. In particular, we have $F(0,...,0,s)=0$ for all $s\in S$. Then, up to shrinking $W$, we have a natural decomposition
@@
More precisely, let $\mathcal{I}$ be a coherent ideal sheaf on $S$ whose cosupport is proper closed subset. Up to shrinking $X$, we assume that $\mathcal{I}$ is globally generated. Let $h\colon W^{\prime}\to W$ be the blowup at the ideal $(x_{1},...,x_{N},\mathcal{I})$. Since the cosupport of $\mathcal{I}$ is strictly contained in $S$, we can denote by $X^{\prime}$ and $S^{\prime}$ the strict transforms of $X$ and $S$ respectively in $W^{\prime}$. Let $\mathcal{J}_{X^{\prime}}$ be the ideal sheaf of $X^{\prime}$ inside $W^{\prime}$.
Then for every point $o^{\prime}\in S^{\prime}$ lying over $o$, there is a neighborhood $U^{\prime}\subseteq W^{\prime}$ of $o^{\prime}$, such that $U^{\prime}$ is isomorphic to an neighborhood of $\mathbf{0}_{N}\times(S^{\prime}\cap U^{\prime})$ in $\mathbb{C}^{N}\times(S^{\prime}\cap U^{\prime})$, and that $h|_{U^{\prime}}\colon U^{\prime}\to W$ can be written in coordinates as
$(x^{\prime}_{1},...,x^{\prime}_{N},s^{\prime})\to(tx^{\prime}_{1},...,tx^{\prime}_{N},h(s^{\prime}))=(x_{1},...,x_{N},s),$
where $t$ is a holomorphic function on $S^{\prime}\cap U$, which is a generator of the invertible ideal sheaf $(h|_{S^{\prime}})^{*}\mathcal{I}$ on $S^{\prime}\cap U^{\prime}$. Furthermore, regarded as a function on $U^{\prime}$, i.e. $t(x^{\prime}_{1},...,x^{\prime}_{N},s^{\prime})=t(s^{\prime})$, $t$ is a generator of the invertible ideal sheaf $h^{*}(x_{1},...,x_{N},\mathcal{I})$ on $U^{\prime}$. Indeed, we can let $U^{\prime}\subseteq W^{\prime}\setminus h^{-1}_{*}H$ and $t=(h|_{S^{\prime}})^{*}\xi$, where $\xi$ belongs to some generating set of $\mathcal{I}$, and $H\subseteq W$ is the zero locus of the function $\xi$ on $W$ defined by $\xi(x_{1},...,x_{N},s)=\xi(s)$.
@@ 3.C. Blowups at a Cartier divisor
Let $f\colon W^{\prime}\to W$ be the blowup at the ideal $\mathcal{I}=(x,y,z,T_{i})$, where $i\in\{1,...,n\}$. Let $X^{\prime}$ and $S^{\prime}$ be the strict transforms of $X$ and $S$ respectively in $W^{\prime}$. Our interest is a defining function of $X^{\prime}$ in a neighborhood of $S^{\prime}$. We note that $f$ induces an isomorphism from $S^{\prime}$ to $S$. There is an open subset $U^{\prime}\subseteq W^{\prime}\setminus f^{-1}_{*}(D_{i})$, where $D_{i}$ is the divisor in $W$ defined by $T_{i}=0$, such that $U^{\prime}\cong\mathbb{D}^{3}\times S$, and that $f|_{U^{\prime}}$ can be written with coordinates as
$(x^{\prime},y^{\prime},z^{\prime},s)\to(T_{i}x,T_{i}y,T_{i}z,s).$
@@
More precisely, let $f\colon W^{\prime}\to W$ be the composition of the following sequence
$W^{\prime}=W^{e}\to W^{e-1}\to\cdots\to W^{0}=W,$
such that each $\varphi^{k}\colon W^{k+1}\to W^{k}$ is the blowup of $W^{k}$ at the center $H_{i}^{k}$ for some $i\in\{1,...,n\}$, where $H_{i}^{k}$ is the preimage of $H_{i}$ in $S^{k}$, and $S^{k}$ is the strict transform of $S$ in $W^{k}$. We say that there are $m$ blowups whose centers are $H_{i}$ if there are $\varphi^{k_{1}},...,\varphi^{k_{m}}$ in the previous sequence, such that $\varphi^{k_{l}}$ is the blowup of $W^{k_{l}}$ at the center $H_{i}^{k_{l}}$ for all $l=1,...,m$. Let $X^{\prime}$ and $S^{\prime}$ be the strict transforms of $X$ and $S$ respectively in $W^{\prime}$. Then $f$ induces an isomorphism from $S^{\prime}$ to $S$. Let $T^{\prime}=\prod_{i=1}^{n}T_{i}^{m_{i}}$, where $m_{i}$ is the number of blowups whose centers are $H_{i}$. Then there is an open neighborhood $U^{\prime}$ of $S^{\prime}$ in $W^{\prime}$, such that $U^{\prime}\cong\mathbb{D}^{3}\times S^{\prime}$ and that $f|_{U^{\prime}}$ can be written with coordinates as
$(x^{\prime},y^{\prime},z^{\prime},s)\mapsto(T^{\prime}x,T^{\prime}y,T^{\prime}z,s).$ (3.1)
@@
By choosing $m_{i}$ greater than the vanishing order of $a(s)b(s)$ along $H_{i}$ for all $i=1,...,n$, the function $a(s)b(s)$ divides $T^{\prime}$. As a result, the defining equation of $X$ can be written as

### v6

@@ 3.A. Functorial resolution of singularities
Let $X$ be a complex analytic variety. Then $X$ admits a functorial (or canonical) desingularization $r\colon\widetilde{X}\rightarrow X$. More precisely, $\widetilde{X}$ is a smooth complex analytic variety, $r$ is a projective bimeromorphic morphism, which is a composition of blowups at smooth centers contained in $X_{\operatorname{\textsubscript{\rm sing}}}$, the $r$-exceptional locus is a snc divisor, and the construction of $r$ commutes with local analytic isomorphisms.
@@
Let $X$ be a complex analytic variety, and let $\mathcal{I}$ be a coherent ideal sheaf on $X$. Then $\mathcal{I}$ admits a functorial (or canonical) principalization $r\colon\widetilde{X}\rightarrow X$. More precisely, $\widetilde{X}$ is a smooth complex analytic variety, $r$ is a projective bimeromorphic morphism, which is a composition of blowups at smooth centers contained in the cosupport of $\mathcal{I}$, $r^{-1}\mathcal{I}\cdot\mathcal{O}_{\widetilde{X}}$ is an invertible sheaf which defines a divisor of snc support, and the construction of $r$ commutes with local analytic isomorphisms.
We will need the following two results on partial desingularizations, which allow us to work locally around components of the singular locus. For example, we assume that $X$ is a complex analytic variety, and that $S_{1}$ and $S_{2}$ are disjoint compact irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$. Let $X_{1}$ and $X_{2}$ be disjoint open subsets of $X$, which contain $S_{1}$ and $S_{2}$ respectively. Suppose that, for $i=1,2$, we have a proper bimeromorphic modification $f_{i}\colon Y_{i}\rightarrow X_{i}$, which is an isomorphism over the smooth locus of $X_{i}$. We remark that there can be an obstruction to glue the modifications $f_{i}$ into a modification on $X$. Indeed, there may be a component $S_{3}$ of $X_{\operatorname{\textsubscript{\rm sing}}}$ which intersects both $S_{1}$ and $S_{2}$. If the modification of $f_{1}$ at general points of $S_{3}\cap X_{1}$ is different from the one of $f_{2}$ at general points of $S_{3}\cap X_{2}$, then we cannot glue $f_{1}$ and $f_{2}$. With the following two results, we can first obtain a modification $\rho\colon X^{\prime}\rightarrow X$ so that the singular loci of $\rho^{-1}(X_{1})$ and $\rho^{-1}(X_{2})$ are compact subsets. Then the modifications on $\rho^{-1}(X_{1})$ and $\rho^{-1}(X_{2})$, which are isomorphic over the smooth loci, are independent and we can always glue them by using Lemma 2.3.
@@
Let $X$ be a normal complex analytic variety and let $S$ be the union of some irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$. Let $\Delta$ be a reduced divisor in $X$, which does not contain any component of $S$. Then there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$ satisfying the following properties. Let $U=X\setminus S$ and $V=f^{-1}(U)$.
@@
Let $r\colon Z\rightarrow X$ be a log resolution of $(X,\Delta)$ such that it can be decomposed into a sequence
$Z=Z_{m}\rightarrow\cdots\rightarrow Z_{0}=X,$
where each $Z_{j+1}\rightarrow Z_{j}$ is the blowup at a smooth irreducible center $C_{j}$ contained in the singular locus (i.e. the non-snc locus) of $(X_{j},\Delta_{j})$, and $\Delta_{j}$ is the sum of the strict transform of $\Delta$ in $X_{j}$ and the whole reduced exceptional divisor of $X_{j}\rightarrow X$. We will construct by induction a sequence of projective bimeromorphic morphisms
$Y=Y_{m}\rightarrow\cdots\rightarrow Y_{0}=X$
such that the following properties hold. The composition $f\colon Y\rightarrow X$ satisfies the properties of the proposition. In particular, we can define $S_{j}$ the strict transform of $S$ in $Y_{j}$, and $Y_{j+1}\rightarrow Y_{j}$ is an isomorphism over general points of every irreducible component of $S_{j}$. Let $V_{j}$ be the preimage of $U$ in $Y_{j}$ and let $W_{j}$ be the preimage of $U$ in $Z_{j}$. Then $V_{j}\cong W_{j}$ under the natural bimeromorphic map $\psi_{j}\colon Z_{j}\dashrightarrow Y_{j}$ over $X$.
Assume that we have constructed $Y_{j}$. If $C_{j}\cap W_{j}=\emptyset$, then we let $Y_{j+1}\rightarrow Y_{j}$ be the identity map. We note that $W_{j+1}\cong W_{j}$. Hence $V_{j+1}\cong W_{j+1}$ under $\psi_{j+1}$.
If $C_{j}\cap W_{j}\neq\emptyset$, then we can define the strict transform $D_{j}$ of $C_{j}$ via $\psi_{j}$. Let $Y_{j+1}\rightarrow Y_{j}$ the blowup of $D_{j}$. It follows that $V_{j+1}\cong W_{j+1}$ under $\psi_{j+1}$. It remains to prove that $Y_{j+1}\rightarrow Y_{j}$ is an isomorphism over general points of every irreducible component of $S_{j}$. It is equivalent to show that $D_{j}$ does not contain any of these irreducible components. Assume by contradiction that $D_{j}$ contains a component $F_{j}$ of $S_{j}$. Then the image $D$ of $D_{j}$ in $X$ contains the one $F$ of $F_{j}$. We note that $D\cap U\neq\emptyset$. On the one hand, by assumption, $F$ is an irreducible component of $S$. On the other hand, $D$ is also the image of $C_{j}$ in $X$. By assumption on $Z$, $D$ is contained in singular locus of $(X,\Delta)$. It follows that there is an irreducible component $P$ of the singular locus of $(X,\Delta)$ such that
@@
In conclusion, $Y_{j+1}\rightarrow Y_{j}$ is an isomorphism over general points of every irreducible component of $S_{j}$. This completes the proof of the proposition. ∎
@@
With the notation in Proposition 3.3, we assume that $X$ has quotient singularities around $S\setminus\Delta$. Let $g\colon X^{\prime}\rightarrow Y$ be the normalization of the blowup of $Y$ at the ideal sheaf $\mathcal{I}$ of $E+f^{-1}_{*}\Delta$, and let $\rho\colon X^{\prime}\rightarrow X$ be the natural morphism. Then the following properties hold.
@@ 3.B. Blowups at an ideal sheaf
Let $X$ be a complex analytic space and let $S\subseteq X$ be a closed subspace. If $g\colon S^{\prime}\rightarrow S$ is a blowup at a coherent ideal sheaf $\mathcal{I}$, then there is a natural blowup $f\colon X^{\prime}\rightarrow X$, such that the restriction of $f$ on the strict transform of $S$ in $X^{\prime}$ is isomorphic to $g$. Indeed, the closed analytic subspace $Z\subseteq S$ defined by $\mathcal{I}$ is naturally a closed analytic subspace of $X$. If we denote the ideal sheaf of $Z$ in $X$ by $\mathcal{I}_{1}$, then we set $f\colon X^{\prime}\rightarrow X$ as the blowup at the ideal sheaf $\mathcal{I}_{1}$. As a consequence, if $g\colon S^{\prime}\rightarrow S$ is now a composition of a sequence of functorial blowups, then there is a natural functorial morphism $f\colon X^{\prime}\rightarrow X$ such that the restriction of $f$ on the strict transform of $S$ in $X^{\prime}$ is isomorphic to $g$.
For an effective Cartier divisor $H\subseteq S$, the blowup of $S$ at (the ideal sheaf of) $H$ is just an isomorphism. However, we can blowup $X$ at $H$ and get a non trivial morphism $f\colon X^{\prime}\rightarrow X$. In this case, $f$ induces an isomorphism from the strict transform $S^{\prime}$ of $S$ to $S$. We will focus on this operation in the next subsection.
In the remainder of this subsection, we let $o\in S\subseteq X$ be a point. We will shrink $S$ and $X$ around $o$ freely. Let $W$ be an open neighborhood of $\{\mathbf{0}_{N}\}\times S$ in $\mathbb{C}^{N}\times S$. We assume that there is a closed immersion from $X$ to $W$ such that the composite inclusion $S\subseteq X\subseteq W$ coincides with $\{\mathbf{0}_{N}\}\times S\subseteq W$. We denote by $\mathcal{J}_{X}$ the ideal sheaf of $X$ inside $W$. Let $F(x_{1},...,x_{N},s)$ be a holomorphic function on $W$, where $(x_{1},...,x_{N})$ are the coordinates of $\mathbb{C}^{N}$ and $s\in S$ is a point. We assume that $F\in\mathcal{J}_{X}$. In particular, we have F(0,,0,s)=0F(0,...,0,s)=0 for all $s\in S$. Then, up to shrinking $W$, we have a natural decomposition
@@
More precisely, let $\mathcal{I}$ be a coherent ideal sheaf on $S$ whose cosupport is proper closed subset. Up to shrinking $X$, we assume that $\mathcal{I}$ is globally generated. Let $h\colon W^{\prime}\rightarrow W$ be the blowup at the ideal $(x_{1},...,x_{N},\mathcal{I})$. Since the cosupport of $\mathcal{I}$ is strictly contained in $S$, we can denote by $X^{\prime}$ and $S^{\prime}$ the strict transforms of $X$ and $S$ respectively in $W^{\prime}$. Let $\mathcal{J}_{X^{\prime}}$ be the ideal sheaf of $X^{\prime}$ inside $W^{\prime}$.
Then for every point $o^{\prime}\in S^{\prime}$ lying over $o$, there is a neighborhood $U^{\prime}\subseteq W^{\prime}$ of $o^{\prime}$, such that $U^{\prime}$ is isomorphic to an neighborhood of $\mathbf{0}_{N}\times(S^{\prime}\cap U^{\prime})$ in $\mathbb{C}^{N}\times(S^{\prime}\cap U^{\prime})$, and that $h|_{U^{\prime}}\colon U^{\prime}\rightarrow W$ can be written in coordinates as
$(x^{\prime}_{1},...,x^{\prime}_{N},s^{\prime})\rightarrow(tx^{\prime}_{1},...,tx^{\prime}_{N},h(s^{\prime}))=(x_{1},...,x_{N},s),$
where $t$ is a holomorphic function on $S^{\prime}\cap U$, which is a generator of the invertible ideal sheaf $(h|_{S^{\prime}})^{*}\mathcal{I}$ on $S^{\prime}\cap U^{\prime}$. Furthermore, regarded as a function on $U^{\prime}$, i.e. $(x^{\prime}_{1},...,x^{\prime}_{N},s^{\prime})=t(s^{\prime}), tt$ is a generator of the invertible ideal sheaf $h^{*}(x_{1},...,x_{N},\mathcal{I})$ on $U^{\prime}$. Indeed, we can let $U^{\prime}\subseteq W^{\prime}\setminus h^{-1}_{*}H$ and $t=(h|_{S^{\prime}})^{*}\xi$, where $\xi$ belongs to some generating set of $\mathcal{I}$, and $H\subseteq W$ is the zero locus of the function $\xi$ on $W$ defined by $\xi(x_{1},...,x_{N},s)=\xi(s)$.
@@ 3.C. Blowups at a Cartier divisor
Let $f\colon W^{\prime}\rightarrow W$ be the blowup at the ideal $\mathcal{I}=(x,y,z,T_{i})$, where $i\in\{1,...,n\}$. Let $X^{\prime}$ and $S^{\prime}$ be the strict transforms of $X$ and $S$ respectively in $W^{\prime}$. Our interest is a defining function of $X^{\prime}$ in a neighborhood of $S^{\prime}$. We note that $f$ induces an isomorphism from $S^{\prime}$ to $S$. There is an open subset $U^{\prime}\subseteq W^{\prime}\setminus f^{-1}_{*}(D_{i})$, where $D_{i}$ is the divisor in $W$ defined by $T_{i}=0$, such that $U^{\prime}\cong\mathbb{D}^{3}\times S$, and that $f|_{U^{\prime}}$ can be written with coordinates as
$(x^{\prime},y^{\prime},z^{\prime},s)\rightarrow(T_{i}x,T_{i}y,T_{i}z,s).$
@@
More precisely, let $f\colon W^{\prime}\rightarrow W$ be the composition of the following sequence
$W^{\prime}=W^{e}\rightarrow W^{e-1}\rightarrow\cdots\rightarrow W^{0}=W,$
such that each $\varphi^{k}\colon W^{k+1}\rightarrow W^{k}$ is the blowup of $W^{k}$ at the center $H_{i}^{k}$ for some $i\in\{1,...,n\}$, where $H_{i}^{k}$ is the preimage of $H_{i}$ in $S^{k}$, and $S^{k}$ is the strict transform of $S$ in $W^{k}$. We say that there are $m$ blowups whose centers are $H_{i}$ if there are $\varphi^{k_{1}},...,\varphi^{k_{m}}$ in the previous sequence, such that $\varphi^{k_{l}}$ is the blowup of $W^{k_{l}}$ at the center $H_{i}^{k_{l}}$ for all l=1,,ml=1,...,m. Let $X^{\prime}$ and $S^{\prime}$ be the strict transforms of $X$ and $S$ respectively in $W^{\prime}$. Then $f$ induces an isomorphism from $S^{\prime}$ to $S$. Let $T^{\prime}=\prod_{i=1}^{n}T_{i}^{m_{i}}$, where $m_{i}$ is the number of blowups whose centers are $H_{i}$. Then there is an open neighborhood $U^{\prime}$ of $S^{\prime}$ in $W^{\prime}$, such that $U^{\prime}\cong\mathbb{D}^{3}\times S^{\prime}$ and that $f|_{U^{\prime}}$ can be written with coordinates as
$(x^{\prime},y^{\prime},z^{\prime},s)\mapstochar\rightarrow(T^{\prime}x,T^{\prime}y,T^{\prime}z,s).$ (3.1)
@@
By choosing $m_{i}$ greater than the vanishing order of $a(s)b(s)$ along $H_{i}$ for all i=1,,ni=1,...,n, the function $a(s)b(s)$ divides $T^{\prime}$. As a result, the defining equation of $X$ can be written as